OpenStudy (anonymous):

Emma has a rope that is 180 cm long. She wants to cut the rope into 45 cm long pieces. How many pieces can Emma make?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

well since she has a rope that size and is simply cuttting it; i believe that all it takes is simple division 180/45 which knowing that 45 + 45 = 90 and 90 + 90 = 180 so; 45 + 45 + 45 + 45 = 180 so she can make up too four pieces since 180/45 = 4

OpenStudy (anonymous):

there was meant to be a square root sign on the 180 and the 45

OpenStudy (anonymous):

it would still be the same, since they both have the square root sign; and the question is asking about the pieces, not a calculation, so if she was cutting she would still end up with four pieces.